What is the probability found by experimenting?
It is experimental or empirical probability.
It is experimental or empirical probability.
It is experimental or empirical probability.
It is experimental or empirical probability.
What is the probability that you rolling factors of 3?
With a single roll of a fair die, the probability is 1/3.
What is the probability of getting at most one tail when two coins are tossed together?
pr(at most one T) = 1 - pr(not two tails) = 1 - 1/2*1/2 = 1 - 1/4 = 3/4
What is joint probability function?
The joint probability function for two variables is a probability function whose domain is a subset of two dimensional space. The joint probability function for discrete random variables X and Y is given as
pr(x, y) = pr(X = x and Y = y). If X and Y are independent random variables then this will equal pr(X =x)*pr(Y = y).
For continuous variables, the joint funtion is defined analogously:
f(x, y) = pr(X < x and Y < y).
What are complementary events in probability?
Complementary events are events that are the complete opposite. The compliment of event A is everything that is not event A.
For example, the complementary event of flipping heads on a coin would be flipping tails.
The complementary event of rolling a 1 or a 2 on a six-sided die would be rolling a 3, 4, 5, or 6.
(The probability of A compliment is equal to 1 minus the probability of A.)
How is probability related to the area under the normal curve?
The Normal curve is a graph of the probability density function of the standard normal distribution and, as is the case with any continuous random variable (RV), the probability that the RV takes a value in a given range is given by the integral of the function between the two limits. In other words, it is the area under the curve between those two values.
What is the probability of getting sum 8 if you throw two dice?
You can get 8 with (2, 6), (3, 5), (4, 4), (5, 3) or (6, 2), each with a probability of (1/6)*(1/6) = 1/36
The the total probability is 5/36 = 0.1388... recurring.
What is the outcome of picking a month of the year and tossing a coin?
There are 24 possible outcomes:
January-Heads,
January-Tails,
February-Heads,
February-Tails,
March-Heads,
and so on.
It depends on the events.
The answer is 0.5*(Total number of events - number of events with probability = 0.5)
That is, discount all events such that their probability (and that of their complement) is exactly a half.
Then half the remaining events will have probabilities that are greater than their complement's.
Samantha Rolls died in 1996, in Vancouver, British Columbia, Canada of heroin overdose.
What is the probability of picking a spade out of a 52 card deck of cards?
One in four.
There are 52 cards in the deck; 13 cards in each of four suits. Spades is one of the four suits, so you have one chance in 4 of picking a spade from a shuffled deck.
What is the probability of the spinner not landing on 1 or 4?
The answer depends on the number of sides on the spinner and what numbers are on it.
What is the probability of something ever happening?
Something just happened. As something would have to happen to make nothing happen, the probability of something else happening is approximately 100%.
If you have a more specific definition of something, probability can be calculated by [# of events] divided by [# of possible outcomes].
For example:
> The chances of rolling a 1 or a 2 on a 6-sided die
How many 7 of clubs in a deck of cards?
In a normal 52-card "Poker" deck, there is one "7 of Clubs" card. In fact, there's only one of every type except the Joker, and there are generally two Joker cards.
Of course, in Pinochle deck, there aren't any sevens at all.
What is the probability of rolling a number greater than 18?
The answer depends on what you are rolling: three or more ordinary dice, or fewer dice with non-standard numbers on them, or a die with some other shape.
What is the probability that N will not have Huntington disease?
0, since letters of the alphabet do not have diseases.
What is the probability that a couple with four children have 2 boys and 2 girls?
1 out of 4. Regardless of what the first two children are, there is a 50/50 chance that each of the following two kids will fulfill the remaining two conditions
What is the probability of 4 boys in a family with 4 children?
In a family with four children, the probability of having four boys is 1 in 16.
What is the probability of not rolling a sum of 5 with 2 dice?
It is 1 - Pr(rolling a sum of 5 with two dice)
= 1 - Pr(1+4 or 2+3 or 3+2 or 4+1) = 1 - 4/36 = 8/9
1. What I Would Do Is Add 2+3+5+6 Which Is 16
2. So Now You Know There Are 16 Marbles
3.Look At How Many Red And Blue Marbles There Are
4. Red=6
5.Blue=2
6. So 6+2=8
7. Your Answer Is 8/16 Or 1/2
8. So The Answer Is 8/16 Or In Simplest Form 1/2